OpenStudy (luigi0210):

Find the area between the curves:      y=2x+2 y=3x^2+1

OpenStudy (anonymous):

Integrate. You can do it.

OpenStudy (anonymous):

Do you know the limits of integration? I suspect if you do limit from -infinity to infinity the whole integral blows up.

OpenStudy (luigi0210):

I forgot how to set up the integral.. They intercept at 1 and -1/3

OpenStudy (anonymous):

The integrand will be the greater function minus the lesser function.

OpenStudy (anonymous):

Looks like \(2x+2\) is the greater function. between the intercepts.

OpenStudy (anonymous):

Yes. So integrate. You don't need really to worry about greater minus lesser. Just take the absolute value of them all.

OpenStudy (luigi0210):

I got a graph:

OpenStudy (anonymous):

So really, what is stopping you?

OpenStudy (anonymous):

You know how to integrate? Integrate A=\[\int\limits_{-1/3}^{1} ((2x+2)-(3x^2+1)) dx\]

OpenStudy (anonymous):

\[ A= \int\limits_{-\frac 13}^1(2x+2)-(3x^2+1)\;dx \]

OpenStudy (luigi0210):

Okay so i end up with: \[\LARGE (x^2+2x)-(x^3+x)\] Right?

OpenStudy (luigi0210):

With the limits of -1/3 to 1 ofc

OpenStudy (anonymous):

Finish it.

OpenStudy (luigi0210):

Finishing it up I get: \[\LARGE -x^3+x^2+x\] \[\LARGE -(1)^3+1^2+1=1\] \[\LARGE -(-\frac{1}{3})^3+(-\frac{1}{3})^2+(-\frac{1}{3})\] So the final answer will be from \[\LARGE (1)-[ -(-\frac{1}{3})^3+(-\frac{1}{3})^2+(-\frac{1}{3})]\] I think..

OpenStudy (anonymous):

That can be simplified dude.

OpenStudy (dan815):